# Extended mass and spheroidal deformation effects on epicyclic frequencies and periapsis shift in quasi-circular orbits

### v2

@@ Abstract
We investigate the effects of extended mass and spheroidal deformation on the periapsis shift of quasi-circular orbits inside a gravitating mass distribution in the Newtonian framework. Focusing on the internal gravitational potential of a spheroidal body with both homogeneous and inhomogeneous density profiles, we elucidate how the ratio of local density to average density governs the extended mass effect on the periapsis shift. By analyzing the orbital angular frequency, along with the radial and vertical epicyclic frequencies, we demonstrate that in the uniform density case (i.e., the Maclaurin spheroid), where the potential takes the form of a harmonic oscillator, the periapsis exhibits a constant retrograde shift of $-\pi$. In contrast, in regions where density inhomogeneity and spheroidal deformation (in both prolate and oblate forms) are significant, the periapsis shift varies with the guiding orbital radius due to local density contrast and deformation effects. The results indicate that oblate deformation suppresses the extended mass effect associated with the ratio of local density to average density, whereas prolate deformation amplifies it. Furthermore, by varying the density distribution parameters, we establish the conditions for orbital stability and identify the emergence of marginally stable orbits.
I Introduction
@@
Recent advances in observations of the Galactic center have provided detailed insights into the gravitational field surrounding the supermassive object [1, 2]. For example, precise tracking of S2’s (also known as S0-2) orbit confirms a subtle prograde periapsis shift that is consistent with the predictions of general relativity [3, 4, 5]. Furthermore, higher-order relativistic cross-term effects in the perihelion precession of Mercury may be detectable by the BepiColombo mission [6, 7].
Motivated by these findings, theoretical research in this field has grown increasingly active. Various approaches have been developed to address periapsis shifts, including the post-Newtonian approximation [8, 9, 10, 11, 12], the numerical relativity [13, 14], and series expansion methods in gravitational theory [15, 16, 18, 17, 19]. Moreover, a novel expansion method has recently been introduced [20, 21]. Since quasi-circular orbits constitute the leading order in the eccentricity expansion, a thorough understanding of them is essential for accurately determining higher-order corrections and periapsis shifts.
@@
For a uniform-density sphere, a test particle moving through only part of the interior experiences a retrograde periapsis shift [22], whereas an orbit entirely within the sphere—reflecting a harmonic oscillator potential—results in a retrograde shift of $-\pi$ (see, e.g., Ref. [23]). Inhomogeneous density distributions further reveal a competition between general relativistic corrections and Newtonian extended mass effects [24]. Furthermore, models incorporating relativistic corrections to the extended mass effects have been applied to scenarios including black holes with matter fields [25, 26], dark matter cores [27, 28], black holes with cloud systems [29, 30, 31], boson stars [32], wormhole [33, 34, 35], and naked singularities [36, 37, 38, 39, 40].
@@
Recently, a general formula for periapsis shifts in static, spherically symmetric spacetimes was proposed, providing a unified framework that transcends previous case studies [41]. In this approach, the formula is expressed in terms of the invariant local curvature, the guiding orbital radius, and the quasi-local mass [see Eq. (2)], which renders it applicable to a variety of gravitational theories, including general relativity. Since curvature directly reflects the energy density and pressure of matter fields through the field equations, this formula is essential for interpreting extended mass effects and deepening our understanding of the phenomenon.
@@
Realistic astrophysical mass distributions frequently deviate from perfect spherical symmetry, thereby significantly affecting the orbital dynamics. [^1] For example, the subtle deformation of the Sun and its impact on the periapsis shift in the solar system have been extensively analyzed using the post-Newtonian method [11]. Recent studies have investigated epicyclic oscillations in Newtonian gravitational fields generated by ellipsoidal mass distributions. In the static, axisymmetric vacuum field of an oblate spheroid with uniform density (i.e., a Maclaurin spheroid), the epicyclic frequencies have been studied in detail [42, 43], revealing a striking similarity to those calculated for the Kerr spacetime in Ref. [44]. In this setting, when the oblateness exceeds a critical value, the radial epicyclic frequency of quasi-circular orbits on the equatorial plane reaches a maximum at a specific radius and subsequently vanishes near the spheroid’s surface, thereby inducing orbital instability. Such behaviors arise from contributions of higher-order mass multipoles [45, 46]. Moreover, analyses have been extended to ellipsoidal models with inhomogeneous density, further including prolate configurations: in oblate models with a relatively gradual central density profile, an unstable region for quasi-circular orbits emerges near the surface (similar to uniform models), whereas oblate models with a steeper central density concentration stabilize the orbits; in prolate models, the radial epicyclic frequency increases monotonically as the guiding orbital radius decreases, thereby ensuring stability over the entire region [47].
@@
Motivated by this background, we examine the periapsis shift in quasi-circular orbits within Newtonian theory by employing a generalized mass distribution model that extends the conventional spherically symmetric case by incorporating a one-parameter deformation. We investigate how this deformation influences extended mass effects by establishing a quantitative relation between the periapsis shift angle and both local and quasi-local average mass densities.
This paper is organized as follows. In Sec. II, we review extended mass effects on the periapsis shift, focusing on the local-to-average density ratio in a static, spherically symmetric model. In Sec. III, we derive the gravitational potential for an ellipsoidal mass distribution, focusing on a solid spheroid. In Sec. IV, we formulate the motion of a test particle on quasi-circular orbits within a reflection symmetric plane, deriving the orbital angular and epicyclic frequencies, and defining the periapsis shift angle. In Sec.V, we analyze the frequency and periapsis shift properties for various density distributions and discuss the effects of source deformation. Finally, in Sec. VI, we summarize our results and outline directions for future research.
II extended mass effects on the periapsis shift of quasi-circular orbits in a spherically symmetric Newtonian system
@@
Recently, a general formula for the periapsis shift angle of quasi-circular orbits of a test particle in a static and spherically symmetric spacetime was presented in Ref. [41] as
$\Delta\phi_{\mathrm{p}}=2\pi\left\{\left[\>\!1-\frac{6m(r)}{r}+\bigg{(}1-\frac{2m(r)}{r}\bigg{)}\frac{\rho(r)+P(r)+2\Pi(r)}{\frac{m(r)}{4\pi r^{3}}+P(r)}+4\pi r^{2}\left[\>\!\rho(r)-3P(r)\>\!\right]\>\!\right]^{-1/2}-1\right\},$ (2)
@@ II.1 Diffuse regime: $\rho/\bar{\rho}\ll 1$
II.2 Homogeneous regime: $\rho/\bar{\rho}\approx 1$
@@ II.3 Spike regime: $\rho/\bar{\rho}\gg 1$
III Gravitational potential of an inhomogeneous spheroid
In this section, we review the gravitational potential inside an inhomogeneous spheroid [50]. Let $x_{i}$ ($i=1,2,3$) denote the Cartesian coordinates in the three-dimensional Euclidean space $\mathbb{R}^{3}$. We consider a solid ellipsoid with semi-axis lengths $a_{1}$, $a_{2}$, and $a_{3}$. We assume that the solid ellipsoid is foliated by thin, homogeneous homeoids, each with semi-axis lengths given by $ma_{i}$ ($i=1,2,3$), where the parameter $m$ varies from 00 at the center to $1$ at the surface. Each homeoid is defined by
@@ III Gravitational potential of an inhomogeneous spheroid
This result implies that the density of this ellipsoid depends solely on $m^{2}$, i.e., $\rho=\rho(m^{2})$. Such an inhomogeneous ellipsoidal mass distribution yields an interior gravitational potential (see Theorem 12 in Ref. [50]),
@@
where $\rho_{0}$ is a constant, and $n$ is a positive number. When $n=0$, Eq. (13) reduces to a constant density within the spheroid, corresponding to the Maclaurin spheroid. Substituting Eq. (13) into Eq. (8) and performing the integration, we obtain
@@
$\Phi_{n}(R,z)=-n!\>\!\pi\>\!G\>\!\rho_{0}\>\!a^{2}\sum_{\begin{subarray}{c}\alpha+\beta+\gamma=n+1\\ \alpha,\beta,\gamma\geq 0\end{subarray}}(-1)^{\alpha+\beta}\frac{C_{\alpha,\beta}}{\alpha!\>\!\beta!\>\!\gamma!}\Big{(}\frac{R}{a}\Big{)}^{2\alpha}\Big{(}\frac{z}{b}\Big{)}^{2\beta},$ (17)
@@
IV Epicyclic oscillations and periapsis shift in quasi-circular orbits
@@
which is derived directly from the Poisson equation for a static, axisymmetric gravitational field [43, 48, 49].
@@
V Properties of the epicyclic frequencies and the periapsis shift angle
V.1 $n=0$
@@
This result directly follows from the fact that the potential $\Phi_{0}(R,0)$ assumes a harmonic oscillator form (i.e., it is proportional to $R^{2}$ up to an additive constant). In other words, the result is derived under the conditions $\rho=\bar{\rho}=\rho_{0}$ (corresponding to a homogeneous regime) and $f=0$ in Eq. (37). It is worth noting, however, that in the vacuum region outside the Maclaurin spheroid, $\Delta\phi_{\mathrm{p}}$ is generally positive [43, 47]. Therefore, we can interpret the constant negativity of $\Delta\phi_{\mathrm{p}}$ in the internal region as a consequence of the extended mass effect in the homogeneous regime, without any additional influence from deformation.
@@ V.2 $n=1$
If $R_{\mathrm{ms}}\leq a$, then $R_{\mathrm{ms}}$ corresponds to the radius of a marginally stable circular orbit inside the oblate spheroid. The right panel in Fig. 2 shows that for the oblateness in the range $e_{0}\leq e<1$ (with $e_{0}=0.929955685$ defined by $R_{\mathrm{ms}}|_{e=e_{0}}=a$), the condition $R_{\mathrm{ms}}<a$ holds. This implies that for $0<e<e_{0}$, the squared radial frequency $\omega_{\mathrm{R}}^{2}$ is positive for all $R\leq a$, ensuring stable circular orbits throughout the spheroid. However, for $e_{0}\leq e<1$, $\omega_{\mathrm{R}}^{2}$ remains positive only for $R<R_{\mathrm{ms}}$, becomes zero at $R=R_{\mathrm{ms}}$, and turns negative for $R_{\mathrm{ms}}<R\leq a$.
@@
where the local mass density is given by $\rho(\xi^{2})=\rho_{0}(1-\xi^{2})$ and the average mass density by $\bar{\rho}(\xi^{2})=\rho_{0}(1-\frac{3}{5}\xi^{2})$. The parameter $f$ is given as
@@ V.3 $n=2$
$\Phi_{2,\mathrm{eff}}=\frac{l^{2}}{2R^{2}}-\frac{\pi\>\!G\rho_{0}a^{2}}{3}$ $\bigg{[}\>\!C_{0,0}-3C_{1,0}\Big{(}\frac{R}{a}\Big{)}^{2}-3C_{0,1}\Big{(}\frac{z}{b}\Big{)}^{2}$ (56)
$+3C_{2,0}\Big{(}\frac{R}{a}\Big{)}^{4}+6C_{1,1}\Big{(}\frac{R}{a}\Big{)}^{2}\Big{(}\frac{z}{b}\Big{)}^{2}+3C_{0,2}\Big{(}\frac{z}{b}\Big{)}^{4}$ (57)
$-C_{3,0}\Big{(}\frac{R}{a}\Big{)}^{6}-3C_{2,1}\Big{(}\frac{R}{a}\Big{)}^{4}\Big{(}\frac{z}{b}\Big{)}^{2}-3C_{1,2}\Big{(}\frac{R}{a}\Big{)}^{2}\Big{(}\frac{z}{b}\Big{)}^{4}-C_{0,3}\Big{(}\frac{z}{b}\Big{)}^{6}\>\!\bigg{]},$ (58)
@@
Figure 4 shows the plots of these squared frequencies. Notably, in the oblate case at $e=0.99$, $\omega^{2}_{\mathrm{R}}$ exhibits two zero crossings, implying the existence of two distinct marginally stable circular orbits.
Figure 4: Plots of the squared orbital angular and epicyclic frequencies for $n=2$. The right panel shows a magnified view of the shaded region shown in the left panel.
@@
The behavior of $\Delta\phi_{\mathrm{p}}$ undergoes a significant transition at the critical value $e_{0}=0.988937349$. The left panel of Fig. 6 shows the region in the $R$-$e$ plane where quasi-circular orbits exist (blue-shaded) and where they do not (unshaded). When $e<e_{0}$, quasi-circular orbits exist over the entire range of $R$, whereas for $e_{0}\leq e<1$, a region in $R$ emerges where quasi-circular orbits are absent (i.e., where $\omega^{2}_{\mathrm{R}}<0$). Note that, unlike in the $n=1$ case where the marginally stable circular orbit appears at the spheroid’s surface, at $e=e_{0}$, it occurs at $R/a=0.969442391$. Consequently, for $e$ in the range $[e_{0},1)$, $\Delta\phi_{\mathrm{p}}$ diverges at those radii corresponding to the two marginally stable circular orbits (see the right panel of Fig. 6).
@@
Figure 6: Left panel: Region in the $R$-$e$ plane for $n=2$ in the oblate case, showing where quasi-circular orbits exist (blue-shaded) and where they do not (unshaded). The boundary, depicted by the blue and orange curves, corresponds to the marginally stable circular orbits, with their intersection marking the critical value $e=0.988937$. Right panel: Plots of $\Delta\phi_{\mathrm{p}}$ for several values of $e$ (with $e>0.988937$), each curve exhibiting divergence at two distinct radii.
V.4 $n=3$

### v3

@@ Abstract
We investigate the effects of extended mass and spheroidal deformation on the periapsis shift of quasi-circular orbits inside a gravitating mass distribution in the Newtonian framework. The analysis is restricted to orbits confined to the reflection-symmetric plane of the spheroidal configuration. Focusing on the internal gravitational potential of a spheroidal body with both homogeneous and inhomogeneous density profiles, we elucidate how the ratio of local density to average density governs the extended mass effect on the periapsis shift. By analyzing the orbital angular frequency, along with the radial and vertical epicyclic frequencies, we demonstrate that in the uniform density case (i.e., the Maclaurin spheroid), where the potential takes the form of a harmonic oscillator, the periapsis exhibits a constant retrograde shift of $-\pi$. In contrast, in regions where density inhomogeneity and spheroidal deformation (in both prolate and oblate forms) are significant, the periapsis shift varies with the guiding orbital radius due to local density contrast and deformation effects. The results indicate that oblate deformation suppresses the extended mass effect associated with the ratio of local density to average density, whereas prolate deformation amplifies it. Furthermore, by varying the density distribution parameters, we establish the conditions for orbital stability and identify the emergence of marginally stable orbits.
keywords:
gravitation – celestial mechanics – black hole physics – dense matter – stars: neutron – galaxies: kinematics and dynamics
1 Introduction
@@
Recent advances in observations of the Galactic center have provided detailed insights into the gravitational field surrounding the supermassive object (Ghez et al., 2000; Schödel et al., 2002). For example, precise tracking of S2’s (also known as S0-2) orbit confirms a subtle prograde periapsis shift that is consistent with the predictions of general relativity (GRAVITY Collaboration et al.,, 2018; Do et al., 2019; Saida et al., 2019). Furthermore, higher-order relativistic cross-term effects in the perihelion precession of Mercury may be detectable by the BepiColombo mission (Will, 2018; Benkhoff et al., 2021).
Motivated by these findings, theoretical research in this field has grown increasingly active. Various approaches have been developed to address periapsis shifts, including the post-Newtonian approximation (Damour & Schaefer, 1988; Damour, Jaranowski & Schaefer, 2000; Yamada & Asada, 2012; Poisson & Will, 2014; Iwata & Yoo, 2016), the numerical relativity (Mroue et al., 2010; Le Tiec et al., 2011), and series expansion methods in gravitational theory (Bini et al., 2005; Vogt & Letelier, 2008; Mak, Leung & Harko, 2018; Tucker & Will, 2019; He, Xu & Jia, 2024). Moreover, a novel expansion method has recently been introduced (Walters, 2018; Katsumata et al., 2025). Since quasi-circular orbits constitute the leading order in the eccentricity expansion, a thorough understanding of them is essential for accurately determining higher-order corrections and periapsis shifts.
@@
For a uniform-density sphere, a test particle moving through only part of the interior experiences a retrograde periapsis shift (Jiang & Lin, 1985), whereas an orbit entirely within the sphere—reflecting a harmonic oscillator potential—results in a retrograde shift of $-\pi$ (see, e.g., Binney & Tremaine (2008)). Inhomogeneous density distributions further reveal a competition between general relativistic corrections and Newtonian extended mass effects (Rubilar & Eckart, 2001). Furthermore, models incorporating relativistic corrections to the extended mass effects have been applied to scenarios including black holes with matter fields (Della Monica, de Martino & de Laurentis, 2022; Bambhaniya et al., 2025), dark matter cores (Argüelles et al., 2022; Igata & Takamori, 2022), black holes with cloud systems (Sadeghian, Ferrer & Will, 2013; Igata et al., 2023; Jusufi, 2023), boson stars (Grould et al., 2017), wormhole (Manna et al., 2019; Deligianni et al., 2021; De Falco, De Laurentis & Capozziello, 2021), and naked singularities (Stefanov, Gyulchev & Yazadjiev, 2013; Bambhaniya et al., 2021; Ota, Kobayashi & Nakashi, 2022; Bambhaniya et al., 2024; Turimov et al., 2024).
@@
Recently, a general formula for periapsis shifts in static, spherically symmetric spacetimes was proposed, providing a unified framework that transcends previous case studies (Harada et al., 2023). In this approach, the formula is expressed in terms of the invariant local curvature, the guiding orbital radius, and the quasi-local mass [see Eq. (2)], which renders it applicable to a variety of gravitational theories, including general relativity. Since curvature directly reflects the energy density and pressure of matter fields through the field equations, this formula is essential for interpreting extended mass effects and deepening our understanding of the phenomenon.
@@
Realistic astrophysical mass distributions frequently deviate from perfect spherical symmetry, thereby significantly affecting the orbital dynamics. [^1] For example, the subtle deformation of the Sun and its impact on the periapsis shift in the solar system have been extensively analyzed using the post-Newtonian method (Poisson & Will, 2014). Recent studies have investigated epicyclic oscillations in Newtonian gravitational fields generated by ellipsoidal mass distributions. In the static, axisymmetric vacuum field of an oblate spheroid with uniform density (i.e., a Maclaurin spheroid), the epicyclic frequencies have been studied in detail (Amsterdamski et al., 2002; Kluźniak & Rosińska, 2013), revealing a striking similarity to those calculated for the Kerr spacetime in Schmidt (2002). In this setting, when the oblateness exceeds a critical value, the radial epicyclic frequency of quasi-circular orbits on the equatorial plane reaches a maximum at a specific radius and subsequently vanishes near the spheroid’s surface, thereby inducing orbital instability. Such behaviors arise from contributions of higher-order mass multipoles (Shibata & Sasaki, 1998; Zdunik & Gourgoulhon, 2001). Moreover, analyses have been extended to ellipsoidal models with inhomogeneous density, further including prolate configurations: in oblate models with a relatively gradual central density profile, an unstable region for quasi-circular orbits emerges near the surface (similar to uniform models), whereas oblate models with a steeper central density concentration stabilize the orbits; in prolate models, the radial epicyclic frequency increases monotonically as the guiding orbital radius decreases, thereby ensuring stability over the entire region (Bollimpalli, 2022).
@@
Motivated by this background, we examine the periapsis shift in quasi-circular orbits within Newtonian theory by employing a generalized mass distribution model that extends the conventional spherically symmetric case by incorporating a one-parameter deformation. Throughout this work, we restrict our analysis to quasi-circular orbits that lie in the reflection-symmetric (equatorial) plane of the mass distribution. We investigate how this deformation influences extended mass effects by establishing a quantitative relation between the periapsis shift angle and both local and quasi-local average mass densities.
This paper is organized as follows. In Sec. 2, we review extended mass effects on the periapsis shift, focusing on the local-to-average density ratio in a static, spherically symmetric model. In Sec. 3, we derive the gravitational potential for an ellipsoidal mass distribution, focusing on a solid spheroid. In Sec. 4, we formulate the motion of a test particle on quasi-circular orbits within a reflection symmetric plane, deriving the orbital angular and epicyclic frequencies, and defining the periapsis shift angle. In Sec.5, we analyze the frequency and periapsis shift properties for various density distributions and discuss the effects of source deformation. Finally, in Sec. 6, we summarize our results and outline directions for future research.
2 Extended mass effects on the periapsis shift of quasi-circular orbits in a spherically symmetric Newtonian system
@@
Recently, a general formula for the periapsis shift angle of quasi-circular orbits of a test particle in a static and spherically symmetric spacetime was presented in Harada et al. (2023) as
$\Delta\phi_{\mathrm{p}}={}$ $2\pi\left\{\left[\>\!1-\frac{6m(r)}{r}+4\pi r^{2}\left[\>\!\rho(r)-3P(r)\>\!\right]\right.\right.$
$+\bigg{(}1-\frac{2m(r)}{r}\bigg{)}\frac{\rho(r)+P(r)+2\Pi(r)}{\frac{m(r)}{4\pi r^{3}}+P(r)}\>\!\bigg{]}^{-1/2}-1\bigg{\}},$ (2)
@@ 2.1 Diffuse regime: $\rho/\bar{\rho}\ll 1$
2.2 Homogeneous regime: $\rho/\bar{\rho}\approx 1$
@@ 2.3 Spike regime: $\rho/\bar{\rho}\gg 1$
3 Gravitational potential of an inhomogeneous spheroid
In this section, we review the gravitational potential inside an inhomogeneous spheroid (Chandrasekhar, 1969). Hereafter, we work in the Newtonian framework and retain the gravitational constant $G$ explicitly, in contrast to the use of geometrized units in Sec. 2. Let $x_{i}$ ($i=1,2,3$) denote the Cartesian coordinates in the three-dimensional Euclidean space $\mathbb{R}^{3}$. We consider a solid ellipsoid with semi-axis lengths $a_{1}$, $a_{2}$, and $a_{3}$, where “solid” refers to a mass distribution that fills the volume, not to mechanical rigidity. We assume that the solid ellipsoid is foliated by thin, homogeneous homeoids, each with semi-axis lengths given by $ma_{i}$ ($i=1,2,3$), where the parameter $m$ varies from 00 at the center to $1$ at the surface. Each homeoid is defined by
@@ 3 Gravitational potential of an inhomogeneous spheroid
This result implies that the density of this ellipsoid depends solely on $m^{2}$, i.e., $\rho=\rho(m^{2})$. Such an inhomogeneous ellipsoidal mass distribution yields an interior gravitational potential (see Theorem 12 in Chandrasekhar (1969)),
@@
where $\rho_{0}$ is a constant, and $n$ is a positive number. When $n=0$, Eq. (13) reduces to a constant density within the spheroid, corresponding to the Maclaurin spheroid. This density profile corresponds to the Ferrers model, which is widely used in galactic dynamics to represent ellipsoidal mass distributions such as galactic bars and bulges (Binney & Tremaine, 2008). Although it does not describe a hydrostatic equilibrium configuration as in polytropic models, nor does it share the extended asymptotic tail of the Plummer profile, it captures the essential feature of central mass concentration. The index $n$ controls the steepness of this concentration, allowing for systematic variation of the density profile. We adopt this model primarily for its analytical tractability, which enables exact evaluation of the internal gravitational potential and facilitates investigation of deformation effects. Substituting Eq. (13) into Eq. (8) and performing the integration, we obtain
@@
$\Phi_{n}(R,z)={}$ $-n!\>\!\pi\>\!G\>\!\rho_{0}\>\!a^{2}$
×$\times\sum_{\begin{subarray}{c}\alpha+\beta+\gamma=n+1\\ \alpha,\beta,\gamma\geq 0\end{subarray}}(-1)^{\alpha+\beta}\frac{C_{\alpha,\beta}}{\alpha!\>\!\beta!\>\!\gamma!}\Big{(}\frac{R}{a}\Big{)}^{2\alpha}\Big{(}\frac{z}{b}\Big{)}^{2\beta},$ (17)
@@
4 Epicyclic oscillations and periapsis shift in quasi-circular orbits
@@
which is derived directly from the Poisson equation for a static, axisymmetric gravitational field (Kluźniak & Rosińska, 2013; Vieira, Kluźniak & Abramowicz, 2017; Delgado, 2022).
@@
5 Properties of the epicyclic frequencies and the periapsis shift angle
5.1 $n=0$
@@
This result directly follows from the fact that the potential $\Phi_{0}(R,0)$ assumes a harmonic oscillator form (i.e., it is proportional to $R^{2}$ up to an additive constant). In other words, the result is derived under the conditions $\rho=\bar{\rho}=\rho_{0}$ (corresponding to a homogeneous regime) and $f=0$ in Eq. (37). It is worth noting, however, that in the vacuum region outside the Maclaurin spheroid, $\Delta\phi_{\mathrm{p}}$ is generally positive (Kluźniak & Rosińska, 2013; Bollimpalli, 2022). Therefore, we can interpret the constant negativity of $\Delta\phi_{\mathrm{p}}$ in the internal region as a consequence of the extended mass effect in the homogeneous regime, without any additional influence from deformation.
@@ 5.2 $n=1$
If $R_{\mathrm{ms}}\leq a$, then $R_{\mathrm{ms}}$ corresponds to the radius of a marginally stable circular orbit (MSCO) inside the oblate spheroid. The right panel in Fig. 2 shows that for the oblateness in the range $e_{0}\leq e<1$ (with $e_{0}=0.929955685$ defined by $R_{\mathrm{ms}}|_{e=e_{0}}=a$), the condition $R_{\mathrm{ms}}<a$ holds. This implies that for $0<e<e_{0}$, the squared radial frequency $\omega_{\mathrm{R}}^{2}$ is positive for all $R\leq a$, ensuring stable circular orbits throughout the spheroid. However, for $e_{0}\leq e<1$, $\omega_{\mathrm{R}}^{2}$ remains positive only for $R<R_{\mathrm{ms}}$, becomes zero at $R=R_{\mathrm{ms}}$, and turns negative for $R_{\mathrm{ms}}<R\leq a$.
@@
where the local mass density is given by $\rho(\xi^{2})=\rho_{0}(1-\xi^{2})$ and the average mass density by $\bar{\rho}(\xi^{2})=\rho_{0}(1-\tfrac{3}{5}\xi^{2})$. The parameter $f$ is given as
@@ 5.3 $n=2$
$\Phi_{2,\mathrm{eff}}={}$ $\frac{l^{2}}{2R^{2}}-\frac{\pi\>\!G\rho_{0}a^{2}}{3}\bigg{[}\>\!C_{0,0}-3C_{1,0}\Big{(}\frac{R}{a}\Big{)}^{2}-3C_{0,1}\Big{(}\frac{z}{b}\Big{)}^{2}$
$+3C_{2,0}\Big{(}\frac{R}{a}\Big{)}^{4}+6C_{1,1}\Big{(}\frac{R}{a}\Big{)}^{2}\Big{(}\frac{z}{b}\Big{)}^{2}+3C_{0,2}\Big{(}\frac{z}{b}\Big{)}^{4}$
$-C_{3,0}\Big{(}\frac{R}{a}\Big{)}^{6}-3C_{2,1}\Big{(}\frac{R}{a}\Big{)}^{4}\Big{(}\frac{z}{b}\Big{)}^{2}$
$-3C_{1,2}\Big{(}\frac{R}{a}\Big{)}^{2}\Big{(}\frac{z}{b}\Big{)}^{4}-C_{0,3}\Big{(}\frac{z}{b}\Big{)}^{6}\>\!\bigg{]},$ (56)
@@
Figure 4 shows the plots of these squared frequencies. Notably, in the oblate case at $e=0.99$, $\omega^{2}_{\mathrm{R}}$ exhibits two zero crossings, implying the existence of two distinct MSCOs.
Figure 4: Plots of the squared orbital angular and epicyclic frequencies for $n=2$. The right panel shows a magnified view of the shaded region shown in the left panel; the two red dots mark the radii where $\omega_{\mathrm{R}}^{2}=0$ for the oblate sequence with $e=0.99$, corresponding to the MSCOs.
@@
The behavior of $\Delta\phi_{\mathrm{p}}$ undergoes a significant transition at the critical value $e_{0}=0.988937349$. The left panel of Fig. 6 shows the region in the $R$-$e$ plane where quasi-circular orbits exist (blue-shaded) and where they do not (unshaded). When $e<e_{0}$, quasi-circular orbits exist over the entire range of $R$, whereas for $e_{0}\leq e<1$, a region in $R$ emerges where quasi-circular orbits are absent (i.e., where $\omega^{2}_{\mathrm{R}}<0$). Note that, unlike in the $n=1$ case where the MSCO appears at the spheroid’s surface, at $e=e_{0}$, it occurs at $R/a=0.969442391$. Consequently, for $e$ in the range $[e_{0},1)$, $\Delta\phi_{\mathrm{p}}$ diverges at those radii corresponding to the two MSCOs (see the right panel of Fig. 6).
Figure 7 shows $\Phi_{2,\mathrm{eff}}(R,0)$ for oblateness $e=0.99$. Two filled‐circle sequences identify distinct families of stable circular orbits: the left-hand sequence forms the inner branch and the right-hand sequence the outer branch. As $l^{2}$ increases, the guiding radius on the inner branch increases monotonically until the branch terminates at the inner MSCO, marked by the left-hand open diamond. Once slightly displaced from this MSCO, the particle eventually crosses the spheroidal surface. Conversely, as $l^{2}$ decreases, the guiding radius on the outer branch decreases monotonically until the branch terminates at the outer MSCO, marked by the right-hand open diamond. A slight displacement from the outer MSCO drives the particle inward, confining it between the outer MSCO and the inner centrifugal barrier.
The radial interval between the two MSCOs contains only a sequence of unstable circular orbits (filled triangles), for which $\omega_{\mathrm{R}}^{2}<0$; hence no epicyclic oscillations can occur there. A particle placed on one of these unstable circular orbits and given a small radial perturbation inevitably departs: if its specific energy exceeds the surface value $\Phi_{2,\mathrm{eff}}(R=a,0)$ it escapes to the exterior, whereas for lower energy if it is reflected by the central or surface barrier and remains trapped between them.
@@
Figure 6: Left panel: Region in the $R$-$e$ plane for $n=2$ in the oblate case, showing where quasi-circular orbits exist (blue-shaded) and where they do not (unshaded). The boundary, depicted by the blue and orange curves, corresponds to the MSCOs, with their intersection marking the critical value $e=0.988937$. Right panel: Plots of $\Delta\phi_{\mathrm{p}}$ for several values of $e$ (with $e>0.988937$), each curve exhibiting divergence at two distinct radii.
Figure 7: Main panel: Equatorial effective potential $\Phi_{2,\mathrm{eff}}(R,0)$ for an inhomogeneous spheroid with $n=2$ and oblateness $e=0.99$, plotted for six values of $\hat{l}^{2}\equiv l^{2}/(\pi G\rho_{0}a^{4})$. Inset: The same curves after subtracting, curve-by-curve, the fixed reference value $\Phi_{2,\mathrm{eff}}(R_{\ast},0)$ with $R_{*}=0.93\,a$. Filled circles mark the locally stable minima, filled triangles indicate the locally unstable maxima, and open diamonds mark the radii of the inner and outer MSCOs.
5.4 $n=3$